# Aggregating Local and Global Features via Selective State Spaces Model for Efficient Image Deblurring

### v1

@@ Abstract
Image deblurring is a process of restoring a high quality image from the corresponding blurred image. Significant progress in this field has been made possible by the emergence of various effective deep learning models, including CNNs and Transformers. However, these methods often face the dilemma between eliminating long-range blur degradation perturbations and maintaining computational efficiency, which hinders their practical application. To address this issue, we propose an efficient image deblurring network that leverages selective structured state spaces model to aggregate enriched and accurate features. Specifically, we design an aggregate local and global block (ALGBlock) to capture and fuse both local invariant properties and non-local information. The ALGBlock consists of two blocks: (1) The local block models local connectivity using simplified channel attention. (2) The global block captures long-range dependency features with linear complexity through selective structured state spaces. Nevertheless, we note that the image details are local features of images, we accentuate the local part for restoration by recalibrating the weight when aggregating the two branches for recovery. Experimental results demonstrate that the proposed method outperforms state-of-the-art approaches on widely used benchmarks, highlighting its superior performance.
@@
Image deblurring, selective state space, features aggregation
@@ I Introduction
Structured state space models [21, 22, 23], notably the enhanced version Mamba, have recently emerged as efficient frameworks due to their ability to capture long-range dependencies with linear complexity. However, Mamba’s [21] recursive image sequence processing method tends to neglect local pixels, while the abundance of hidden states in the state space equation often results in channel redundancy, thereby impeding channel feature learning. Given the critical importance of local and channel features in image deblurring, directly applying the state space model often leads to poor performance.
Taking into account the above analyses, a natural question arises: Is it feasible to design a network that efficiently aggregates local and global features for image deblurring? To achieve this objective, we propose ALGNet, with several key components. Firstly, we develop an efficient global block that utilizes selective structured state spaces to capture long-range dependency information essential for high-quality image deblurring. Secondly, to address the issues of local pixel forgetting and channel redundancy in the state space equation, we design a local block to model local connectivity using simplified channel attention. Thirdly, as image details primarily consist of local features, we emphasize the importance of the local block in restoration by recalibrating the weights through the learnable factor when aggregating the two blocks for recovery. Finally, to alleviate training difficulty, we implement multiple scales of input and output modes. As illustrated in Figure 1, our ALGNet model achieves state-of-the-art performance while preserving computational efficiency compared to existing methods.
@@
    We propose an efficiently aggregates local and global features network (ALGNet) that leverages selective structured state spaces model for image deblurring.
@@
    We design an aggregate local and global block (ALGBlock) consisting of both global and local components. The global block utilizes selective structured state spaces model to capture long-range dependency information. And the local block models local connectivity using simplified channel attention and address the issues of local pixel forgetting and channel redundancy in the state space equation.
@@
    We present an aggregate methods that emphasize the importance of the local block in restoration by recalibrating the weights through the learnable factor.
@@ II Related Work
Due to the content-dependent global receptive field, the transformer architecture [27] has recently gained much popularity in image restoration [28, 15, 16, 17, 18, 29, 30], demonstrating superior performance compared to previous CNN-based baselines. IPT [28] employs a Transformer-based multi-head multi-tail architecture, proposing a pre-trained model for image restoration tasks. However, image deblurring often deals with high-resolution images, and the attention mechanism in Transformers incurs quadratic time complexity, resulting in significant computational overhead. In order to reduce the computational cost, Uformer [18], SwinIR [30] and U2former [15] computes self-attention based on a window. Nonetheless, the window-based approach still falls short in fully exploring the information within each patch. Restormer [16] and MRLPFNet [19] compute self-attention across channels rather than the spatial dimension, resulting in the linear complexity. However, this approach fails to fully exploit the spatial information. FFTformer [14] explores the property of the frequency domain to estimate the scaled dot-product attention, but need corresponding inverse Fourier transform, leading to additional computation overhead.
II-D State Space Models.
State space models [31, 32, 21, 33, 22, 23] have recently emerged as efficient frameworks due to their ability to capture long-range dependencies with linear complexity. S4 [31] is the first structured SSM to model long-range dependency. S5 [32] propose the diagonal SSM approximation and computed recurrently with the parallel scan. Mega [33] introduced a simplification of S4 [31] to be real- instead of complex- valued, giving it an interpretation of being an exponential moving average. SGConv [34] and LongConv [35] focus on the convolutional representation of S4 and create global or long convolution kernels with different parameterizations. Mamba [21] propose a selective mechanism and hardware-aware parallel algorithm. Many vision tasks start to employ Mamba to tackle image classification [36, 37], image segmentation [38], and so on.
In this work, we leverage selective structured state spaces model to capture long-range dependency information. In order to address the issues of local pixel forgetting and channel redundancy in the state space equation, we model local connectivity using simplified channel attention.
@@ III Method
In this section, we first outline the overall pipeline of our ALGNet. Subsequently, we delve into the details of the proposed ALGBlock which contains global block and local block, and features aggregation (FA).
Figure 2: Overall architecture of ALGNet.
@@ III-A Overall Pipeline
The overall pipeline of our proposed ALGNet, shown in Figure 2, adopts a single U-shaped architecture for image restoration.
Given a degraded image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, ALGNet initially applies a convolution to acquire shallow features $\mathbf{F_{0}}\in\mathbb{R}^{H\times W\times C}$ ($H,W,C$ are the feature map height, width, and channel number, respectively). These shallow features undergo a four-scale encoder sub-network, progressively decreasing resolution while expanding channels. It’s essential to note the use of multi-input and multi-output mechanisms for improved training. The low-resolution degraded images are incorporated into the main path through the Convs and concatenation, followed by convolution to adjust channels. The in-depth features then enter a middle block, and the resulting deepest features feed into a four-scale decoder, gradually restoring features to the original size. During this process, the encoder features are concatenated with the decoder features to facilitate the reconstruction. Finally, we refine features to generate residual image $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$ to which degraded image is added to obtain the restored image: $\mathbf{\hat{I}}=\mathbf{X}+\mathbf{I}$. It’s important to note that the three low-resolution results are solely used for training. To facilitate the frequency selection process, we optimize the proposed network with the following loss function:
@@
where $\mathcal{F}$ represents fast Fourier transform, and the parameters $\lambda$ and $\delta$ control the relative importance of loss terms, which are set to $0.1$ and $0.05$, respectively.
III-B Global Block
Long-range dependency information is important for image deblurring, and transformer-based methods [16, 14, 19, 15, 18] are starting to become mainstream due to the limited receptive field of CNNs. However, image deblurring often deals with high-resolution images, and the attention mechanism in Transformers incurs quadratic time complexity, resulting in significant computational overhead. Although some design window-based attention or compute self-attention across channels rather than the spatial dimension, this methods may cause the information loss. To this end, we design a global block with linear complexity to image deblurring, as shown in Figure 2(c). Specifically, given an input feature $F_{in}$, we first reshape and normalization it by layer norm (LN). Then, it will go through two parallel branches. In the first branch, the feature channel is expanded by a linear layer, SiLU activation function, together with the SSM layer. In the second branch, the features channel is also expanded by a linear layer followed by the SiLU activation function. After that, features from the two branches are aggregated with the element-wise multiplication. Finally, the channel number is projected back and reshape to the original size. The total process can be defined as :
$F_{in}^{1}$ $=SiLU(Linear(Reshape(LN(F_{in}))),$ (5)
$F_{in}^{2}$ $=SSM(SiLU(W_{1D}^{1}(Linear(Reshape(LN(F_{in})))))$
$F_{G}$ $=Reshape(Linear(F_{in}^{1}\otimes F_{in}^{2}))$
III-C Local Block
Although the selective state space model [21] has the ability of global effective receptive field as well as linear computational complexity, it can result in the problem of local pixel forgetting and channel redundancy caused by the excessive hidden state number. To address the above challenges, we design a local block to model local connectivity and facilitate the expressive power of different channels. The local block shown in Figure 2(c), it is a simplified channel attention. Given a features $F_{in}$, the local features $F_{L}$ can be obtained by:
$F_{in}^{{}^{\prime}}$ $=F_{in}\otimes W_{p}^{(1)}(GAP(F_{in})),$ (6)
$F_{L}$ $=W_{p}^{(2)}(F_{in}^{{}^{\prime}})$
where $W_{p}^{(\cdot)}$ is the $1\times 1$ convolution, GAP is the global average pool
III-D Features Aggregation (FA)
@@
As image details primarily consist of local features, we design a features aggregation(FA) module, as illustrated in Figure 2(b), to emphasize the importance of the local block in restoration by recalibrating the weights through the learnable factor when aggregating the two blocks for recovery. Specifically, given the global features $F_{G}$ and local features $F_{L}$, the aggregate process can be defined as:
$F_{A}=F_{G}+WF_{L}$ (7)
@@
where $W$ denotes the learnable parameters, which are directly optimized by backpropagation and initialized as 1. It is worth mentioning that our design is extremely lightweight, as it does not introduce additional convolution layers.
@@ IV-A Experimental Settings
In this section, we introduce the details of the used datasets, and training details.
TABLE I: Image motion deblurring results. The best and second best scores are highlighted and underlined. Our ALGNet-B and ALGNet are trained only on the GoPro dataset but achieves the highest and second-highest scores on the average of the effects on both datasets, respectively.
Methods PSNR $\uparrow$ SSIM $\uparrow$
DeblurGAN-v2 [13] 29.55 0.934
SPAIR [39] 32.06 0.953
MIMO-UNet++ [5] 32.45 0.957
MPRNet [6] 32.66 0.959
MPRNet-local [6] 33.31 0.964
HINet [7] 32.71 0.959
HINet-local [7] 33.08 0.962
Uformer [18] 32.97 0.967
MSFS-Net [40] 32.73 0.959
MSFS-Net-local [40] 33.46 0.964
NAFNet-32 [4] 32.83 0.960
NAFNet-64 [4] 33.62 0.967
Restormer [16] 32.92 0.961
Restormer-local [16] 33.57 0.966
Stripformer [17] 33.08 0.962
IRNeXt [3] 33.16 0.962
SFNet [25] 33.27 0.963
FSNet [26] 33.29 0.963
ALGNet(Ours) 33.49 0.964
ALGNet-B(Ours) 34.05 0.969
TABLE II: Image motion deblurring results.
Methods PSNR $\uparrow$ SSIM $\uparrow$
DeblurGAN [11] $24.51$ $0.871$
DeblurGAN-v2 [13] 26.61 0.875
DBGAN [12] 28.94 0.915
MT-RNN [41] 29.15 0.918
DMPHN [42] 29.09 0.924
SPAIR [39] 30.29 0.931
MIMO-UNet++ [5] 29.99 0.930
MPRNet [6] 30.96 0.939
MPRNet-local [6] 31.19 0.945
Uformer [18] 30.83 0.952
MSFS-Net [40] 31.05 0.941
MSFS-Net-local [40] 31.30 0.943
Restormer [16] 31.22 0.942
Restormer-local [16] 31.49 0.945
Stripformer [17] 31.03 0.940
SFNet [25] 31.10 0.941
FSNet [26] 31.05 0.941
ALGNet(Ours) 31.68 0.957

### v2

# Learning Enriched Features via Selective State Spaces Model for Efficient Image Deblurring

@@ Abstract
Image deblurring aims to restore a high-quality image from its corresponding blurred. The emergence of CNNs and Transformers has enabled significant progress. However, these methods often face the dilemma between eliminating long-range degradation perturbations and maintaining computational efficiency. While the selective state space model (SSM) shows promise in modeling long-range dependencies with linear complexity, it also encounters challenges such as local pixel forgetting and channel redundancy. To address this issue, we propose an efficient image deblurring network that leverages selective state spaces model to aggregate enriched and accurate features. Specifically, we introduce an aggregate local and global information block (ALGBlock) designed to effectively capture and integrate both local invariant properties and non-local information. The ALGBlock comprises two primary modules: a module for capturing local and global features (CLGF), and a feature aggregation module (FA). The CLGF module is composed of two branches: the global branch captures long-range dependency features via a selective state spaces model, while the local branch employs simplified channel attention to model local connectivity, thereby reducing local pixel forgetting and channel redundancy. In addition, we design a FA module to accentuate the local part by recalibrating the weight during the aggregation of the two branches for restoration. Experimental results demonstrate that the proposed method outperforms state-of-the-art approaches on widely used benchmarks.
@@
Image deblurring, state spaces model, features aggregation
@@ I Introduction
State space models [21, 22, 23], notably the enhanced version Mamba, have recently emerged as efficient frameworks due to their ability to capture long-range dependencies with linear complexity. However, Mamba’s [21] recursive image sequence processing method tends to neglect local pixels, while the abundance of hidden states in the state space equation often results in channel redundancy, thereby impeding channel feature learning. Given the critical importance of local and channel features in image deblurring, directly applying the state space model often leads to poor performance. MambaIR [24] introduces the vision state space module, which utilizes a four-direction unfolding strategy to scan along four different directions for local enhancement, and incorporates channel attention to mitigate channel redundancy. Nevertheless, this four-directional scanning approach and the computation of state spaces result in increased computational overhead, potentially sacrificing the advantage of low computational resource utilization offered by the state spaces equation.
Taking into account the above analyses, a natural question arises: Is it feasible to design a network that efficiently aggregates local and global features for image deblurring? To achieve this objective, we propose ALGNet, with several key components. Specifically, we present an aggregate local and global information block (ALGBlock) aimed at efficiently capturing and merging both local invariant properties and long-range dependencies. This ALGBlock consists of two key modules: a module dedicated to capturing local and global features (CLGF), and a feature aggregation module (FA). The CLGF module is further divided into two branches: the global branch, which utilizes a selective state spaces model to capture long-range dependency features with linear complexity, and the local branch, which incorporates simplified channel attention to effectively model local connectivity. This combination not only addresses issues like local pixel forgetting and channel redundancy but also empowers the network to capture more enriched and precise features. Additionally, given that image details are predominantly comprised of local features of images [25], we design the FA module to underscores the significance of the local information in the restoration process by dynamically recalibrating the weights through a learnable factor during the aggregation of the CLGF two branches for restoration. Finally, we implement multiple scales for both input and output modes, aiming to alleviate training difficulty. As illustrated in Figure 1, our ALGNet model achieves state-of-the-art performance while preserving computational efficiency compared to existing methods.
@@
    We propose ALGNet, an efficient network for aggregating enriched and precise features leveraging a selective state spaces model for image deblurring. ALGNet consists of multiple ALGBlocks, each with a capturing local and global features module (CLGF) and a feature aggregation module (FA).
@@
    We design the CLGF module to capture long-range dependency features using a selective state spaces model, while employing simplified channel attention to model local connectivity, thus reducing local pixel forgetting and channel redundancy.
@@
    We present the FA module to emphasize the importance of the local features in restoration by recalibrating the weights through the learnable factor.
@@ II Related Work
[…]
II-D State Spaces Model.
State spaces models [33, 34, 21, 35, 22, 23] have recently emerged as efficient frameworks due to their ability to capture long-range dependencies with linear complexity. S4 [33] is the first structured SSM to model long-range dependency. S5 [34] propose the diagonal SSM approximation and computed recurrently with the parallel scan. Mega [35] introduced a simplification of S4 [33] to be real- instead of complex- valued, giving it an interpretation of being an exponential moving average. SGConv [36] and LongConv [37] focus on the convolutional representation of S4 and create global or long convolution kernels with different parameterizations. Mamba [21] propose a selective mechanism and hardware-aware parallel algorithm. Many vision tasks start to employ Mamba to tackle image classification [38, 39], image segmentation [40], and so on. However, the standard Mamba model still encounters issues with local pixel forgetting and channel redundancy when applied to image restoration tasks. To tackle this challenge, MambaIR [24] adopts a four-direction unfolding strategy to scan along four different directions and integrates channel attention. Nevertheless, this four-directional scanning approach results in increased computational overhead. In this work, we design an ALGBlock to efficiently capture and integrate both local invariant properties and long-range dependencies through a selective state spaces model with lower computational cost. As depicted in Figure 1, our ALGNet outperforms MambaIR [24] while reducing computational costs by 96.1%.
Figure 2: Overall architecture of ALGNet. (a) ALGNet consists of several ALGBlocks and adopts the multi-input and multi-output strategies for image restoration. (b) ALGBlock comprises two primary modules: a module for capturing local and global features (CLGF), and a feature aggregation module (FA). The CLGF module is composed of two branches: (c) the local branch to model local connectivity, while (d) the global branch captures long-range dependency features.
@@ III Method
In this section, we first outline the overall pipeline of our ALGNet. Subsequently, we delve into the details of the proposed ALGBlock, which includes the capturing local and global features module (CLGF) and the feature aggregation module (FA).
@@ III-A Overall Pipeline
The overall pipeline of our proposed ALGNet, shown in Figure 2, adopts a single U-shaped architecture for image deblurring. Given a degraded image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, ALGNet initially applies a convolution to acquire shallow features $\mathbf{F_{0}}\in\mathbb{R}^{H\times W\times C}$ ($H,W,C$ are the feature map height, width, and channel number, respectively). These shallow features undergo a four-scale encoder sub-network, progressively decreasing resolution while expanding channels. It’s essential to note the use of multi-input and multi-output mechanisms for improved training. The low-resolution degraded images are incorporated into the main path through the Convs (consists of multiple convolutions and ReLU) and concatenation, followed by convolution to adjust channels. The in-depth features then enter a middle block, and the resulting deepest features feed into a four-scale decoder, gradually restoring features to the original size. During this process, the encoder features are concatenated with the decoder features to facilitate the reconstruction. Finally, we refine features to generate residual image $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$ to which degraded image is added to obtain the restored image: $\mathbf{\hat{I}}=\mathbf{X}+\mathbf{I}$. It’s important to note that the three low-resolution results are solely used for training.
We optimize the proposed network ALGNet with the following loss function:
@@
where $\mathcal{F}$ represents fast Fourier transform, and the parameters $\lambda$ and $\delta$ control the relative importance of loss terms, which are set to $0.1$ and $0.05$ as in [6, 28], respectively.
III-B Capturing Local and Global Features Module (CLGF)
Transformer-based models [16, 14, 19, 15, 18] address the limitations of CNNs, such as a limited receptive field and lack of adaptability to input content. They excel in modeling non-local information, leading to high-quality image reconstruction, and have emerged as the dominant method for image deblurring. However, image deblurring commonly involves processing high-resolution images, and the attention mechanism in Transformers introduces quadratic time complexity, leading to considerable computational overhead. While it’s possible to mitigate computational consumption by utilizing window-based attention [18, 15] or channel-wise attention [16, 19], these methods inevitably lead to information loss.
To address this challenge, we design the capturing local and global features module (CLGF) depicted in Figure 2(b), aiming to capture long-range dependency features and model local connectivity with linear complexity. Specifically, given an input tensor $X_{l-1}$, we initially process it through Layer Normalization (LN), Convolution, and Simple Gate (SG) to obtain spatial features $X^{s}_{l-1}$ as follows:
$X^{s}_{l-1}$ $=SG(f_{3\times 3}^{dwc}(f_{1\times 1}^{c}(LN(X_{l-1}))))$ (5)
$SG(X_{f0})$ $=X_{f1}\otimes X_{f2}$
where $f_{3\times 3}^{dwc}$ denotes the $3\times 3$ depth-wise convolution, $f_{1\times 1}^{c}$ represents $1\times 1$ convolution. $SG(\cdot)$ is the simple gate, employed as a replacement for the nonlinear activation function. For a given input $X_{f0}\in\ \mathbb{R}^{H\times W\times C}$, SG initially splits it into two features $X_{f1}\in\ \mathbb{R}^{H\times W\times\frac{C}{2}}$ and $X_{f2}\in\ \mathbb{R}^{H\times W\times\frac{C}{2}}$ along channel dimension. Subsequently, SG calculates the $X_{f1},X_{f2}$ using a linear gate. Next, we feed the spatial features $X^{s}_{l-1}$ through both the global branch and the local branch to capture global features $F_{G}$ and local features $F_{L}$, respectively.
In the global branch, depicted in Figure 2(c), we opt for a state-space model (SSM) instead of Transformers to capture long-distance dependencies, ensuring linear complexity. Specifically, starting with an input feature $X^{s}_{l-1}$, we first reshape and normalize it using layer normalization (LN). Subsequently, it undergoes processing through two parallel branches. In the top branch, the feature channels are expanded by a linear layer, followed by activation through the SiLU function. In the bottom branch, the feature channels are expanded by a linear layer followed by the SiLU activation function, along with the selective state spaces model layer. The SSM inspired by the particular continuous system that maps a 1-dimensional function or sequence $x(t)\in\mathbb{R}\rightarrow y(t)\in\mathbb{R}$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$ as follows:
$h^{{}^{\prime}}(t)$ $=\textbf{A}h(t)+\textbf{B}x(t)$ (6)
$y(t)$ $=\textbf{C}h(t)$
where $\mathbf{A}\in\mathbb{R}^{N\times N},\mathbf{B}\in\mathbb{R}^{N\times 1},\mathbf{C}\in\mathbb{R}^{1\times N}$ are four parameters, and $N$ is the state size. SSM first transform the continuous parameters A, B to discrete parameters $\mathbf{\overline{A},\overline{B}}$ through fixed formulas $\mathbf{\overline{A}}=exp(\square\mathbf{A})$ and $\mathbf{\overline{B}}=(\square\mathbf{A})^{-1}exp(\square\mathbf{A}-I)\cdot\square\mathbf{B}$, where $\square$ denotes the timescale parameter. After the discretization, the model can be computed as a linear recurrence way:
$h_{t}$ $=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t}$ (7)
$y_{t}$ $=\textbf{C}h_{t}$
or a global convolution way:
$\mathbf{\overline{K}}$ $=(\mathbf{C\overline{B},C\overline{AB},...,C\overline{A}^{k-1}\overline{B}})$ (8)
[…]
After that, features from the two branches are aggregated with the element-wise multiplication. Finally, the channel number is projected back and reshape to the original size. The total process can be defined as:
$F_{t}$ $=SiLU(Linear(Reshape(LN(X^{s}_{l-1}))),$ (9)
$F_{b}$ $=SSM(SiLU(f_{1\times 1}^{c}(Linear(Reshape(LN(X^{s}_{l-1})))))$
$F_{G}$ $=Reshape(Linear(F_{t}\otimes F_{b}))$
Figure 3: (a) Local pixels (highlighted by the red dashed line) are susceptible to being forgotten in the flattened 1D sequence due to the extensive distance. (b) Following [24], we apply ReLU and global average pooling to the outputs of the global branch to obtain channel activation values. However, a considerable portion of channels remain inactive, indicating channel redundancy.
Despite the selective SSM’s ability to capture long-range dependencies with linear computational complexity, it can result in issues like local pixel forgetting and channel redundancy, primarily due to the flattening strategy and an excessive number of hidden states. As shown in Figure 3(a), when the 2D feature map is flattened into a 1D sequence, adjacent pixels (e.g., sequence number 1 and 11) become widely separated, leading to the issue of pixel forgetting. Additionally, as per [24], we visualize the activation results for different channels in Figure 3(b) and observe significant channel redundancy attributed to the larger number of hidden states in the selective SSM. To address the above challenges, we equip our CLGF with a local branch to model local connectivity and facilitate the expressive power of different channels. The local branch shown in Figure 2(c), it is a simplified channel attention. Given the spatial features $X^{s}_{l-1}$, the local features $F_{L}$ can be obtained by:
$F_{L}^{{}^{\prime}}$ $=X^{s}_{l-1}\otimes f_{1\times 1}^{c}(GAP(X^{s}_{l-1})),$ (10)
$F_{L}$ $=f_{1\times 1}^{c}(F_{L}^{{}^{\prime}})$
where GAP is the global average pool. Noted that, in our CLGF module, we initially capture the spatial feature $X^{s}_{l-1}$, which aggregates neighboring information, before feeding it to the global branch. This approach effectively reduces the problem of local pixel forgetting.
III-C Feature Aggregation Module (FA)
@@
Given that image details primarily consist of local features, we design a feature aggregation (FA) module, depicted in Figure 2(b), to highlight the significance of the local block in restoration. This is achieved by dynamically recalibrating the weights through a learnable factor during the aggregation of the two blocks for recovery. Specifically, given the global features $F_{G}$ and local features $F_{L}$, the aggregate process can be defined as:
$F_{A}=F_{G}\oplus WF_{L}$ (11)
@@
where $W$ represents the learnable parameters, directly optimized by backpropagation and initialized as 1. It’s worth noting that our design is exceptionally lightweight, as it does not introduce additional convolution layers. Finally, to capture richer and more accurate information, we refine the aggregated features $F_{A}$ to obtain the output feature $X_{l}$ of the ALGBlock as follows:
$X_{l}=F_{A}\oplus f_{1\times 1}^{c}(SG(f_{1\times 1}^{c}(LN(X_{l-1}\oplus F_{A}))))$ (12)